Grade 12Physical Chemistry

Find the number of photons emitted per minute by a 25 W source of monochromatic light of wavelenght 5000 ang strom

Frsg
9 Years agoGrade 12

1 Answer

Khimraj
9 Years ago
No of photon is given by equation P*t = nhc/λWhere P is power t is timeh is Planck constantc is velocity of light.n = (25*60*5000*10^-10)/(6.6"10^-34*3*10^8) = 3.79*10^21